OpenStudy (anonymous):

The flower garden has the shape of a right triangle. 68 feet of a perennial border forms the hypotenuse of the triangle, and one leg is 28 feet longer than the other leg. Find the length of the legs

OpenStudy (anonymous):

32 and 60

OpenStudy (anonymous):

hyp = 68, leg = 28+x, leg2 = x, & a^2 = b^2 + c^2 and solve